# Advancing Averaged Primer Vector Theory with Bang-Bang Control and Eclipsing

### v1

@@ 3.2 Multi-arc averaging
Introduced by Tarzi [23], the multi-arc approach to bang-bang optimal control separates the averaging integral into different regions of thrust on/off depending on the roots of the eclipsing and switching function. Figure 3 shows a diagram of a single period with 6 sections. Each section, or “arc”, is treated separately in the averaging integral using the exact boundaries defined by the switching and eclipsing function roots. The segmenting of the averaging integral mitigates the need for a large quadrature number to accurately capture proportions of thrusting and coasting. The arcs with no thrusting due to shadow are treated separately from the arcs that have no thrusting due to the bang-bang optimal control law. Each arc uses fixed values of either 0 or 1 for and , depending on the sign of the eclipsing and switching functions, respectively. The multi-arc averaged dynamics approximate the true bang-bang control of the unaveraged minimum fuel problem.
The M roots of the switching function from Eq. 14 are computed exactly using a new approach. When the switching function is fully expanded in terms of B, a sixth order polynomial of tan can be shown to have the same roots when all variables except are held constant within the averaging integral. The switching function roots are calculated at every integration step when using averaged dynamics, so the use of a simpler “switching polynomial”, , speeds up the overall process. A symbolic manipulator is used to show equivalency between the roots of the switching function in Eq. 14 and those of in Eq. 23:
@@
The form of Eq. 23 implies that a maximum of 6 relevant roots exist, depending on the coefficients, noting that any complex roots are not relevant. Therefore, a maximum number of three separate arcs of thrusting are possible within a single period in the averaging context for a minimum fuel problem. This theoretical result, consistent with the numerical results of [40], implies a maximum number of optimal thrusts per revolution in the low-thrust limit. This new result provides insight on Edelbaum’s question of “how many impulses” are optimal for transfers between orbits [41, 42]. Even in direct optimization, the number of impulses or thrust arcs per revolution has a large impact on run time and resolution of the problem. A theoretical upper bound on the number of thrusts per revolution is therefore a significant result for practical mission design. This theoretical result was independently developed and first presented to the public by the authors in January, 2024 [1]. A similar result was presumably developed simultaneously, and subsequently published by Wu et al. in March 2024 [26].
The eclipsing function is much simpler than the switching function, with at most two roots for a single revolution. The roots are calculated using a Newton-Raphson routine with initial guesses found analytically from a cylindrical shadow model [43]. After the roots of the switching and eclipsing functions are calculated, the averaged Hamiltonian is broken down into a sum across arcs:
@@
The dynamics derivative calculations require the Leibniz integral rule, because the integral limits are now functions of the differentiation variables. In general form, the Leibniz integral rule [44] is
@@
The Leibniz integral rule terms in Eqs. 26 and 27 capture the effect of co-state jumps due to the eclipse constraint in the dynamics, and naturally smooth them out over the averaging period. These terms are functionally equivalent to the Dirac delta terms used by Mazzini[22]. Note also that the Leibniz terms are related to the additional terms (Eq. 29 in Ref. [12]) attributed to sensitivities propagated across bang-bang switches. Figure 4 shows how the quadrature points and Leibniz boundary terms contribute to the co-state dynamics for the same hypothetical multi-arc period. For calculation of these terms, both the derivative and the or are evaluated at true longitude roots of the eclipsing and switching functions. The partial derivative of the switching boundary terms are computed assuming the switching function constraint is satisfied:
For the eclipse entrance/exit points, the situation is simpler, because the eclipsing function is not a function of the co-state:
@@ 3.3 Eclipsing singularity
The new Leibniz boundary terms in Eq. 26 smoothly incorporate the co-state jumps into the averaged equations of motion. As eclipsing arcs get smaller and approach zero length, though, these extra terms increase in magnitude. Unfortunately, they become singular in the limit when the arc disappears. The averaged co-state jumps used in Mazzini’s averaging model [22] similarly become singular. The singularity can be shown mathematically by breaking down a hypothetical averaging period with an eclipse and no coasting arcs:
where are the true longitude values for the entrance and exit of the eclipse, respectively. As an eclipsing arc approaches zero length, term 1 of Eq. 31 approaches infinity, while term 2 stays finite. A trajectory propagated with unaveraged dynamics will only be affected by this singularity if it happens to cross a point where the eclipse arc is very small (approximately less than radians) and incorporates the eclipsing constraint in a discontinuous way [11]. However, a trajectory propagated with averaged dynamics will encounter this singularity if any part of the averaging integral contains a very small eclipsing arc, which occurs nearly every time eclipsing is present. A trajectory propagated with averaged dynamics and a variable step integrator will significantly slow down when approaching the singularity, and will not be able to propagate through it without enforcing a minimum step size. Additionally, a trajectory that is forced to propagate through the singularity will accumulate errors in the equations of motion.
@@
where is the eclipsing arc size in radians. This redefinition causes term 2 of Eq. 31 to approach zero as the eclipsing arcs become small, effectively removing the singularity by canceling with the increase of term 1. A small is preferred, so that the redefinition has a minimal effect on the propagated trajectory. However, a larger with a gradually decreasing function results in a faster propagation. A variable step propagator will always slow down as eclipsing arcs disappear, but the extent of slow-down is affected by this choice. Trade studies are conducted on the function and the arc size where the fix begins, to balance these two effects. Linear through 12th order polynomials are tested with fixed points at and . The following value of and polynomial for are selected to balance compute time and realism:
@@
Wang and Topputo [11] use a similar technique of varying the floor of a discontinuous eclipsing constraint in their approach to the indirect method. Their model is unaveraged, but converges better when the eclipsing constraint floor is slowly lowered using a continuation scheme. The fix shown here is motivated primarily by a singularity in the equations of motion, but improves convergence as well. The overall flow of calculating the multi-arc averaged dynamics is shown in Alg. 1.
@@ 4 Variational Equations
The state transition matrix (STM) of the full augmented state vector is useful for a variety of applications, including solving boundary value problems. Previous works using averaged dynamics either do not mention the STM [20, 22] or use a finite difference method [23]. In this work, the STM, , is integrated along with the combined state and co-state using the variational equations. While tedious to derive, this approach is faster and more accurate. The STM is integrated using the typical variational equations:
where y is the augmented state and co-state vector, and a flipped vector is defined for notational convenience:
The averaged dynamics require extra attention to account for the Leibniz integral terms. The roots of the eclipsing and switching functions, are functions of the state variables, and must be considered in the total derivative. Therefore, the Jacobian of the dynamics is calculated using the quadrature formulation of Eqs. 27 and 26:
5 Results
Before the results are shown, an approximate unaveraged dynamics model is introduced. Discontinuous bang-bang control and eclipsing constraints have been used with unaveraged optimal control by Wang and Topputo [11], Russell [12], and others. This approach typically requires a propagator with integrator events, complicated STM calculations, and explicit co-state jump calculations. A common approach to approximate the unaveraged model uses smooth Heaviside approximations for bang-bang control and the eclipsing constraint, and eliminates the need for event finding or jumps in the co-state STMs. This approximate model is introduced here solely for validation of the averaged model. The approximate Heavisides are set to be extremely steep and as close to true bang-bang control as possible to make the comparison to the averaged dynamics sufficiently accurate. The use of approximate Heavisides for smooth bang-bang control and eclipsing is beyond the current scope. Future work will include approximate Heavisides in detail in both the unaveraged and averaged models. For more details on approximate Heaviside implementations in optimal control, see [14] and [36].
Two examples are presented: A Geostationary Transfer Orbit (GTO) to Geostationary Orbit (GEO) transfer is shown to validate the averaged model through comparison with unaveraged dynamics. The averaged and unaveraged transfers are optimized separately to show how the optimal averaged trajectory can be used as a proxy for the optimal unaveraged trajectory. A Low Earth Orbit (LEO) to GEO optimal transfer is shown using only averaged dynamics. This latter example is almost 500 revolutions and is intractable to practically optimize with unaveraged dynamics.
@@ 5.1 Approximate Unaveraged Dynamics
The approximate unaveraged dynamics make trajectory propagation faster and smoother by substituting smooth Heaviside approximations for bang-bang control and the eclipsing constraint [14]. The optimal control derivation and resulting equations of motion are very similar to those derived in Section 2 with only a few adjustments.
First, the eclipsing constraint, Eq. 5, is changed to a smooth Heaviside approximation:
where is a smoothing parameter. As is reduced to zero, Eq. 38 approximates the discontinuous eclipsing constraint, Eq. 5. As a result of the smoothed eclipsing constraint, the co-state jumps are also smoothed, but will approach discrete jumps as approaches zero. This L2-norm based Heaviside approximation is a an efficient alternative to the trigonometric form [45, 46, 47].
@@
where is a smoothing parameter that is gradually reduced towards zero to approximate the true minimum fuel problem, Eq. 9. The resulting Hamiltonian is:
By applying the smoothing approximation in the cost function definition, the derived optimal control law results in the same L2-norm based form of the Heaviside approximation used in Eq. 38:
Besides the new definitions in Eq. 38 and Eq. 41, all other equations from Section 2 are the same. When compared with averaged dynamics with bang-bang control and eclipsing shown in Sec. 3, the smoothing parameters are set to the order of .
@@
A 48-revolution, optimal minimum-fuel GTO to GEO transfer is presented to compare the averaged and unaveraged dynamics models. The states and co-states of the optimal averaged trajectory are directly compared with the optimal unaveraged trajectory for the first time in the literature. The boundary conditions and spacecraft details are given in Tables 2 and 2, respectively. The Sun position and velocity data are calculated using the SPICE for MATLAB toolkit [48], with a launch date of Apr 1, 2008 (260280065.0 ET). The trajectory is propagated using a Runge-Kutta 8(9) variable step integrator with relative and absolute tolerances set to . The averaged trajectory is propagated using in Eq. 28. The unaveraged trajectory is propagated with smoothing parameters of to approximate bang-bang control and compensate for scaling differences between the switching and eclipsing function amplitudes during the transfer. These values are selected by propagating an initial guess and iterating on the smoothing parameters until the steepness of and look consistent. The full perturbation is included using modified equinoctial elements in the RTN frame [29]:
@@
where the subscripts and represent the initial and final values, respectively, and the transversality conditions are a result of free final mass and true longitude. The initial guess for the co-states of the averaged trajectory was all zeros and the initial guess for the co-states of the unaveraged trajectory was the optimal initial co-states of the averaged trajectory. The optimized initial co-state values for both trajectories are given in the Appendix.
| Parameter | GTO | GEO |
| Semi-Major Axis | 24505 km | 42165 km |
| Eccentricity | 0.725 | 0 |
| Inclination | 28.5 deg | 0 |
| Argument of Periapse | 0 | 0 |
| R.A. of the Asc. Node | 0 | 0 |
| True Anomaly | 0 | free |
@@
| Mass | 100 kg |
@@
| Launch date | 260280065 ET |
| J2 | 0.00108263 |
| Prop Tolerances | |
| Quad Param, | 6 |
The unaveraged trajectory is shown in three dimensions in Fig. 6(a), with thrusting and shadowing details in Fig. 6(b). The averaged trajectory is shown with time information in Fig. 6(c). The sparsity of the points in the averaged trajectory plot show the ability of the variable step integrator to take large steps with averaged dynamics. The variable step integrator takes more frequent steps, though, when the number of thrusting arcs change. The time histories of the states and co-states are compared in Fig. 7. The final mass is a key statistic in the mission design process and a good measure of the ability to use the optimal averaged trajectory as a proxy for the optimal unaveraged trajectory. The optimal unaveraged final mass is 93.638 kilograms and the optimal averaged final mass is 93.645 kilograms. The co-state flattens near , because the trajectory no longer encounters any eclipsing and this co-state is only affected by the presence of eclipsing. The averaged trajectory closely follows the unaveraged trajectory, providing a validation of the averaged dynamics model as an approximation. Some averaged co-states maintain the same time history structure as the unaveraged co-states but become slightly out of sync because they have no transversality conditions in solving the BVP.
The thrust profile is shown in Fig. 8. There is one instance during the transfer where the thrust does not completely switch off (highlighted in the first zoom-in panel), because the optimal switch happens so fast that the smoothing parameters are not small enough to capture such a short arc. In Fig. 9(a), more detail is shown for the first 10 revolutions of state and Fig. 9(b) shows the near-discrete co-state jumps of from eclipsing that the averaged dynamics smoothly incorporates. The number of switches and eclipsing events are tracked in Fig. 11. The cumulative number of integration steps are tracked in Fig. 11, where the unaveraged trajectory requires approximately 60 times more steps. The averaged trajectory takes more frequent steps when the number of thrusting or eclipsing arcs per revolution changes. This slow-down when using a variable step integrator is a inherent downside of averaged dynamics, but typically will only occur a few times during transfers.
The root mean square of the relative difference between the unaveraged and averaged trajectory states and co-states is shown in Fig. 13. This difference is calculated across all states and co-states using:
@@
The variational equations calculated in section 4 are verified with this trajectory. The STM propagated with the variational equations is compared to the STM calculated using the complex step method, a machine-precision, numerical approach to calculating derivatives described in [49] and [50]. The complex step STM is considered truth for comparison purposes. The relative error for each STM entry is calculated by normalizing by the maximum value in the row or column of each entry:
Extra care is taken to ensure that the trajectories propagated with complex step and with the variational equations follow the exact same path. This path uniformity can be achieved by using fixed step propagation, among other methods [50]. The maximum relative STM error for different quadrature parameters, , is shown in Fig. 13. The number of quadrature points per arc does not affect the accuracy of the STM. Therefore, boundary value problems can be solved efficiently, even with low-fidelity dynamics parameters.
@@
A 486-revolution, minimum-fuel optimal GTO to GEO transfer is shown to showcase averaged dynamics. The boundary conditions are identical to Table 2 from the previous example. The spacecraft details are given in Table 3. The optimized initial co-state values are given in the Appendix. The initial guess for the co-states of the averaged trajectory was again all zeros. The averaged trajectory is shown with time information in Fig. 15 and the cumulative number of integration steps are tracked in Fig. 15. Despite the significant increase in the number of revolutions (486 vs. 48), only about 50 percent more integration steps are required compared to the previous transfer. The averaged dynamics eliminate periodic changes in states and co-states, causing the integration steps to scale with qualitative changes like the number of thrusting or eclipsing arcs per revolution instead of the number of total revolutions. The time histories of the states and co-states are shown in Fig. 16 with gray dashed lines for the averaged time histories from the previous example. The structure is marginally different from the previous example despite the same boundary conditions, because the eclipsing seasons are different as a result of a different maximum thrust and time-of-flight.
| Parameter | Value |
| Mass | 100 kg |
| Isp | 3100 sec |
| Max Thrust | 0.01 N |
| Min Thrust | 0 N |
| ToF | 350 Days |
| GM | 398600 km3/sec2 |
| Launch date | 252417665 ET |
| J2 | 0.00108263 |
| Prop Tolerances | |
| Quad Param, | 8 |

### v3

@@ 3.2 Multi-arc averaging
Introduced by Tarzi [tarzi], the multi-arc approach to bang-bang optimal control separates the averaging integral into different regions of thrust on/off depending on the roots of the eclipsing and switching function. Figure 2 shows a diagram of a single period with 6 sections. Each section, or “arc”, is treated separately in the averaging integral using the exact boundaries defined by the switching and eclipsing function roots. The segmenting of the averaging integral mitigates the need for a large quadrature number to accurately capture proportions of thrusting and coasting. The arcs with no thrusting due to shadow are treated separately from the arcs that have no thrusting due to the bang-bang optimal control law. Each arc uses fixed values of either 0 or 1 for and , depending on the sign of the eclipsing and switching functions, respectively. The multi-arc averaged dynamics approximate the true bang-bang control of the unaveraged minimum fuel problem.
The M roots of the switching function from Eq. (16) are computed exactly using a new approach. When the switching function is fully expanded in terms of B, a sixth-order polynomial of tan can be shown to have the same roots when all variables except are held constant within the averaging integral. The switching function roots are calculated at every integration step when using averaged dynamics, so the use of a simpler “switching polynomial”, , speeds up the overall process. A symbolic manipulator is used to show equivalency between the roots of the switching function in Eq. (16) and those of in Eq. (26):
@@
The form of Eq. (26) implies that a maximum of 6 relevant roots exist, depending on the coefficients, noting that any complex roots are not relevant. Therefore, a maximum number of three separate arcs of thrusting are possible within a single period in the averaging context for a minimum fuel problem. This theoretical result, consistent with the numerical results of [taheri2020many], implies a maximum number of optimal thrusts per revolution in the low-thrust limit. Previous works investigating the structure of the switching function have set limits of 8 unique roots per period specifically during coasting in unaveraged minimum fuel problems [Pan2012, Jamison2010] with lower limits when additional constraints are included. The new averaged dynamics-based upper limit of 6 roots provides further insight on Edelbaum’s question of “how many impulses” are optimal for transfers between orbits when averaging assumptions hold [edelbaum_howmanyimpulses, saloglu]. Even in direct optimization, the number of impulses or thrust arcs per revolution has a large impact on run time and resolution of the problem. A theoretical upper bound on the number of thrusts per revolution is therefore a significant result for practical mission design. This theoretical result was independently developed and first presented to the public by the authors in January, 2024 [lifset]. A similar result was presumably developed simultaneously, and subsequently published by Wu et al. in March 2024 [wu_avg].
The eclipsing function is much simpler than the switching function, with at most two roots for a single revolution. The roots are calculated using a Newton-Raphson routine with initial guesses found analytically from a cylindrical shadow model [vallado2001fundamentals]. After the roots of the switching and eclipsing functions are calculated, the averaged Hamiltonian is broken down into a sum across arcs:
@@
The dynamics derivative calculations require the Leibniz integral rule, because the integral limits are now functions of the differentiation variables. In general form, the Leibniz integral rule [Protter1985] is
@@
The Leibniz integral rule terms in Eqs. (29) and (30) capture the effect of costate jumps due to the eclipse constraint in the dynamics, and naturally smooth them out over the averaging period. These terms are functionally equivalent to the Dirac delta terms used by Mazzini[mazzini]. Note also that the Leibniz terms are related to the additional terms attributed to sensitivities propagated across bang-bang switches (see Eq. (29) in Ref. [russell2007primer]). Figure 4 shows how the quadrature points and Leibniz boundary terms contribute to the costate dynamics for the same hypothetical multi-arc period. For calculation of these terms, both the derivative and the or are evaluated at true longitude roots of the eclipsing and switching functions. The partial derivative of the switching boundary terms are computed assuming the switching function constraint is satisfied:
For the eclipse entrance/exit points, the situation is simpler, because the eclipsing function is not a function of the costate:
@@ 3.3 Eclipsing singularity
The new Leibniz boundary terms in Eq. (29) smoothly incorporate the costate jumps into the averaged equations of motion. As eclipsing arcs get smaller and approach zero length, though, these extra terms increase in magnitude. Unfortunately, they become singular in the limit when the arc disappears. The averaged costate jumps used in Mazzini’s averaging model [mazzini2] similarly become singular. The singularity can be shown mathematically by breaking down a hypothetical averaging period with an eclipse and no coasting arcs:
where are the true longitude values for the entrance and exit of the eclipse, respectively. As an eclipsing arc approaches zero length, term 1 of Eq. (34) approaches infinity, while term 2 stays finite. A trajectory propagated with unaveraged dynamics will only be affected by this singularity if it happens to cross a point where the eclipse arc is very small (approximately less than radians) and incorporates the eclipsing constraint in a discontinuous way [wangtopputo]. However, a trajectory propagated with averaged dynamics will encounter this singularity if any part of the averaging integral contains a very small eclipsing arc, which occurs nearly every time eclipsing is present. A trajectory propagated with averaged dynamics and a variable step integrator will significantly slow down when approaching the singularity, and will not be able to propagate through it without enforcing a minimum step size. Additionally, a trajectory that is forced to propagate through the singularity will accumulate errors in the equations of motion.
@@
where is the eclipsing arc size in radians. This redefinition causes term 2 of Eq. (34) to approach zero as the eclipsing arcs become small, effectively removing the singularity by canceling with the increase of term 1. A small is preferred, so that the redefinition has a minimal effect on the propagated trajectory. However, a larger with a gradually decreasing function results in a faster propagation. A variable step propagator will always slow down as eclipsing arcs disappear, but the extent of slow-down is affected by this choice. Trade studies are conducted on the function and the arc size where the fix begins, to balance these two effects. Linear through 12th order polynomials are tested with fixed points at and . The following value of and polynomial for are selected to balance compute time and realism:
@@
Wang and Topputo [wangtopputo] use a similar technique of varying the floor of a discontinuous eclipsing constraint in their approach to the indirect method. Their model is unaveraged, but converges better when the eclipsing constraint floor is slowly lowered using a continuation scheme. The fix shown here is motivated primarily by a singularity in the equations of motion, but improves convergence as well. The overall flow of calculating the multi-arc averaged dynamics is shown in Alg. 1.
@@ 4 Variational Equations
The state transition matrix (STM) of the full augmented state vector is useful for a variety of applications, including solving boundary value problems. Previous works using averaged dynamics either do not mention the STM when solving the BVP [geffroy, mazzini] or use a finite difference method to calculate it [tarzi]. In this work, the STM, , is integrated along with the combined states and costates using the variational equations. While tedious to derive, this approach is faster and more accurate. The variational equations for averaged dynamics are explicitly provided here for the first time in the literature. It is possible that prior authors have utilized the variational equations to compute STMs and solve BVPs, but details of their calculation, including methods to avoid singularities, have not yet appeared in the averaging literature. The STM is integrated using the typical approach with the variational equations:
where y is the augmented state and costate vector, and a flipped vector is defined for notational convenience:
The averaged dynamics require extra attention to account for the Leibniz integral terms. The roots of the eclipsing and switching functions, are functions of the state variables, and must be considered in the total derivative. Therefore, the Jacobian of the dynamics is calculated using the quadrature formulation of Eqs. (30) and (29):
5 Example Optimal Trajectories
Before the results are shown, an approximate unaveraged dynamics model is introduced. Discontinuous bang-bang control and eclipsing constraints have been used with unaveraged optimal control by Wang and Topputo [wangtopputo], Russell [russell2007primer], and others. This approach typically requires a propagator with integrator events, complicated STM calculations, and explicit costate jump calculations. A common approach to approximate the unaveraged model uses smooth Heaviside approximations for bang-bang control and the eclipsing constraint, and eliminates the need for event finding or jumps in the costate STMs. This approximate model is introduced here solely for validation of the averaged model. The approximate Heavisides are set to be extremely steep and as close to true bang-bang control as possible to make the comparison to the averaged dynamics sufficiently accurate. The use of approximate Heavisides for smooth bang-bang control and eclipsing in averaged dynamics is beyond the current scope. For more details on approximate Heaviside implementations in optimal control, see [taheri] and [bertrandepenoy].
Two examples are presented: A GTO to GEO transfer is shown to validate the averaged model through comparison with unaveraged dynamics. The averaged and unaveraged transfers are optimized separately to show how the optimal averaged trajectory can be used as a proxy for the optimal unaveraged trajectory. A GTO to GEO optimal transfer is then shown using only averaged dynamics. This latter example is almost 500 revolutions and is intractable to practically optimize with unaveraged dynamics.
@@ 5.1 Approximate Unaveraged Dynamics
The approximate unaveraged dynamics make trajectory propagation faster and smoother by substituting smooth Heaviside approximations for bang-bang control and the eclipsing constraint [taheri]. The optimal control derivation and resulting equations of motion are very similar to those derived in Section 2 with only a few adjustments.
First, the eclipsing constraint, Eq. (6), is changed to a smooth Heaviside approximation:
where is a smoothing parameter. As is reduced to zero, Eq. (41) approximates the discontinuous eclipsing constraint, Eq. (6). It should be noted that the smooth Heaviside approximations asymptotically approach zero and one, but the averaged model sets the thrust to be equal to exactly zero or one. As a result of the smoothed eclipsing constraint, the costate jumps are also smoothed, but will approach discrete jumps as approaches zero. This L2-norm based Heaviside approximation is a an efficient alternative to the trigonometric form [taheri_L2norm, Mall_L2norm, Heidrich_L2norm].
@@
where is a smoothing parameter that is gradually reduced towards zero to approximate the true minimum fuel problem, Eq. (10). The resulting Hamiltonian is:
By applying the smoothing approximation in the cost function definition, the derived optimal control law results in the same L2-norm based form of the Heaviside approximation used in Eq. (41):
Besides the new definitions in Eq. (41) and Eq. (44), all other equations from Section 2 are the same. When compared with averaged dynamics with bang-bang control and eclipsing shown in Sec. 3, the smoothing parameters are set to the order of . The smooth Heaviside approximations introduced here are unrelated in mathematical structure with the Heaviside redefinition introduced in Sec. 3.3. The two Heaviside adjustments are related in motivation, though, because they both benefit the simulation of trajectories.
@@
A 48-revolution, optimal minimum-fuel GTO to GEO transfer is presented to compare the averaged and unaveraged dynamics models. The states and costates of the optimal averaged trajectory are directly compared with the optimal unaveraged trajectory for the first time in the literature. The transfer parameters and details are given in Tables 5.2 and 5.2, respectively. Note that the parameters of this first example require an unrealistically high power level. However, they are chosen to simply validate the averaged dynamics model. A more practical transfer is shown in the later example. The Sun position and velocity data are calculated using the SPICE for MATLAB toolkit [spice], with a launch date of Apr 1, 2008 (260280065.0 ET). The trajectory is propagated using a Runge-Kutta 8(9) variable step integrator with relative and absolute tolerances set to . The averaged trajectory is propagated using in Eq. (31). The unaveraged trajectory is propagated with smoothing parameters of to approximate bang-bang control and compensate for scaling differences between the switching and eclipsing function amplitudes during the transfer. These values are selected by propagating an initial guess and iterating on the smoothing parameters until the steepness of and look consistent. The full perturbation is included using modified equinoctial elements in the RTN frame [kechichian]:
@@
where the subscripts and represent the initial and final values, respectively, and the transversality conditions are a result of free final mass and true longitude. The initial guess for the costates of the averaged trajectory was all zeros and the initial guess for the costates of the unaveraged trajectory was the optimal initial costates of the averaged trajectory. The optimized initial costate values for both trajectories are given in the Appendix.
@@
| Semimajor axis {GTO, GEO} | {24505 km, 42165 km} |
| Eccentricity {GTO, GEO} | {0.725, 0} |
| Inclination {GTO, GEO} | {28.5 deg, 0 deg} |
| AOP {GTO, GEO} | {0 deg, 0 deg} |
| RAAN {GTO, GEO} | {0 deg, 0 deg} |
| TA {GTO, GEO} | {0 deg, free} |
| Initial Mass | 100 kg |
@@
| Integration tolerance | |
| Parameter | 48 rev transfer | 486 rev transfer |
| Thrust | [0, 0.2] N | [0, 0.01] N |
| ToF | 30 Days | 350 Days |
| Initial epoch | 260280065 ET | 252417665 ET |
| Quad Param, | 6 | 8 |
| Final mass (avg) | 93.645 kg | 91.946 kg |
| Final mass (unavg) | 93.638 kg | - |
| (avg) | 1.996079 km/s | 2.552701 km/s |
| (unavg) | 1.998351 km/s | - |
| Eclipse arcs | 34 | 389 |
| Coast arcs | 59 | 377 |
| Integration steps (avg) | 281 | 385 |
| Integration steps (unavg) | 17,756 | - |
The unaveraged trajectory is shown in three dimensions in Fig. 6(a), with thrusting and shadowing details in Fig. 6(b). The averaged trajectory is shown with time information in Fig. 6(c). The sparsity of the points in the averaged trajectory plot show the ability of the variable step integrator to take large steps with averaged dynamics. The variable step integrator takes more frequent steps, though, when the number of thrusting arcs change. The time histories of the states and costates are compared in Fig. 7. Additionally, the time histories of states are shown as classical Keplerian orbital elements in Fig. 8. Note that the RAAN and AOP become undefined when eccentricity equals zero. The final mass is a key statistic in the mission design process and a good measure of the ability to use the optimal averaged trajectory as a proxy for the optimal unaveraged trajectory. The optimal unaveraged final mass is 93.638 kilograms (corresponding to 1.998351 km/s ) and the optimal averaged final mass is 93.645 kilograms (corresponding to 1.996079 km/s ). The costate flattens near , because the trajectory no longer encounters any eclipsing and this costate is only affected by the presence of eclipsing. The averaged trajectory closely follows the unaveraged trajectory, providing a validation of the averaged dynamics model as an approximation. Some averaged costates maintain the same time history structure as the unaveraged costates but become slightly out of sync because they have no transversality conditions in solving the boundary value problem.
The thrust profile is shown in Fig. 9. There is one instance during the transfer where the thrust does not completely switch off (highlighted in the first zoom-in panel), because the optimal switch happens so fast that the smoothing parameters are not small enough to capture such a short arc. In Fig. 10(a), more detail is shown for the first 10 revolutions of state and Fig. 10(b) shows the near-discrete costate jumps of from eclipsing that the averaged dynamics smoothly incorporates. The number of switches and eclipsing events are tracked in Fig. 12. The cumulative number of integration steps are tracked in Fig. 12, where the unaveraged trajectory requires approximately 60 times more steps. The averaged trajectory takes more frequent steps when the number of thrusting or eclipsing arcs per revolution changes. This slow-down when using a variable step integrator is a inherent downside of averaged dynamics, but typically will only occur a few times during transfers.
The root mean square of the relative difference between the unaveraged and averaged trajectory states and costates is shown in Fig. 14. This difference is calculated across all states and costates using:
@@
The variational equations calculated in section 4 are verified with this trajectory. The STM propagated with the variational equations is compared to the STM calculated using the complex step method, a machine-precision, numerical approach to calculating derivatives described in [martins2003complex] and [pellegrini_STM]. The complex step STM is considered truth for comparison purposes. The relative error for each STM entry is calculated by normalizing by the maximum value in the row or column of each entry:
Extra care is taken to ensure that the trajectories propagated with complex step and with the variational equations follow the exact same path. This path uniformity can be achieved by using fixed step propagation, among other methods [pellegrini_STM]. The maximum relative STM error for different quadrature parameters, , is shown in Fig. 14. The number of quadrature points per arc does not affect the accuracy of the STM. Therefore, boundary value problems can be solved efficiently, even with low-fidelity dynamics parameters.
@@
A 486-revolution, minimum-fuel optimal GTO to GEO transfer is shown to showcase averaged dynamics. The boundary conditions and transfer parameters are identical to Table 5.2 from the previous example. The transfer details are given in Table 5.2. The optimized initial costate values are given in the Appendix. The initial guess for the costates of the averaged trajectory was again all zeros. The averaged trajectory is shown with time information in Fig. 16 and the cumulative number of integration steps are tracked in Fig. 16. Despite the significant increase in the number of revolutions (486 vs. 48), only about 50 percent more integration steps are required compared to the previous transfer. The averaged dynamics eliminate periodic changes in states and costates, causing the integration steps to scale with qualitative changes like the number of thrusting or eclipsing arcs per revolution instead of the number of total revolutions. The time histories of the states and costates are shown in Fig. 17 with gray dashed lines for the averaged time histories from the previous example. The structure is marginally different from the previous example despite the same boundary conditions, because the eclipsing seasons are different as a result of a different maximum thrust and time-of-flight. Additionally, the time histories of the states are shown as classical Keplerian orbital elements in Fig. 18 with gray dashed lines for the averaged Keplerian time histories from the previous example. The final mass is 91.946 kg, corresponding to a 2.552701 km/s .